OpenStudy (anonymous):

which number can be multiplied on both sides to make x=12? 3/4x=9. A)4 B)4/3 C)1/4 D) -3/4

OpenStudy (jango_in_dtown):

B) 4/3 since (4/3)(3/4)x=(4/3)9 implis x=4.3=12

OpenStudy (jango_in_dtown):

@allineedisyourforever

OpenStudy (anonymous):

Can you help me with 1 more? @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

ok

OpenStudy (anonymous):

Solve. -1/5x=6 A)-6/5 B)6 1/5 C)-1 D)-30

OpenStudy (jango_in_dtown):

-1/5 x=6 multiply both sides by-5 and you will get x=-30 so option D is correct